OpenStudy (anonymous):

there is a pop quiz with five questions- 4 true/false and one multiple choice. each true/false question has two possible answers and the multiple choice question has four possible answers(a-d). Whats the probability that the student, guessing randomly, answers all five questions correctly?

OpenStudy (tkhunny):

You can't "guess" "randomly". It doesn't work. p(1 of 2) = 1/2 p(1 of 4) = 1/4